# Practical Aspects on Solving Differential Equations Using Deep Learning: A Primer

### v3

@@ 15 L_init = ((y0 - y_ic)**2)
Figure 10: First-order ordinary differential equation solution. A Solution of Equation (19) in the interval $[0,1]$. The blue solid line shows the analytic solution $y(t)=2\exp(-t)$, and the orange crosses represent the solution obtained by the Deep Galerkin (DG) method. The approximated solution is close to the analytic one (mean absolute error MAE$=0.0017$). B Evolution of loss for the DG method.
@@ 6 Fredholm Integral Equations
[…]
where #$\#[a,b]$ is the length of $[a,b]$, which is $\frac{pi}{2}$ for equation (26), and the number of samples (random points) we draw from a uniform distribution $\mathcal{U}(a,b)=\mathcal{U}(0,\frac{\pi}{2})$ is $k=50$. So for every point $j\in k$, we draw a random sample, we pass it to the neural network $\mathcal{F}$ to obtain the $\hat{y}$ and then we plug that in equation (28) to get the integral term $\mathcal{I}$. After we evaluate the integral term, we proceed in minimizing the loss function (27). The code listing 10 shows the Pytorch implementation of the loss function for this problem, where in lines $8$-$12$, we evaluate the integral part described by Equation (28), and in line $15$, we compute the solution $\hat{y}$ using the DGM-like neural network. Finally, in line $17$, we estimate the loss function based on Equation (27).
[…]
@@
We use the Adam optimizer with its default parameters and learning rate $\eta=0.0001$. We train the neural network for $3,000$ iterations. In each iteration we uniformly draw $n=32$ (batch size) random numbers from $[0,\frac{\pi}{2}]$ representing the independent variable $x$ and we feed the mini-batch to the neural network.
@@
Figure 12 A shows the analytic solution of equation (26) (blue line) and the approximated solution evaluated with the trained neural network on the interval $[0,\frac{\pi}{2}]$ with $50$ grid points. We see that the approximated solution overlays with the analytic one, and the MAE=$0.0134$. Panel 12 B shows the loss during training. We observe how the loss drops until convergence after about $500$ iterations.

### v4

@@ 15 L_init = ((y0 - y_ic)**2)
Figure 10: First-order ordinary differential equation solution. A Solution of Equation (19) in the interval $[0,1]$. The blue solid line shows the analytic solution $y(t)=2\exp(-t)$, and the orange crosses represent the solution obtained by the Deep Galerkin (DG) method. The approximated solution is close to the analytic one (mean absolute error MAE=0.0017=0.0017). B Evolution of loss for the DG method.
@@ 6 Fredholm Integral Equations
$\mathcal{I}$ $=\int_{a}^{b}K(x,t)y(t)dt\approx\frac{\#[a,b]}{k}\sum_{j=1}^{k}K(x_{i},t_{j})\hat{y}(t_{j})=\frac{\#[0,\frac{\pi}{2}]}{50}\sum_{j=1}^{50}\sin(x_{i})\cos(t_{j})\hat{y}(t_{j})$ (28)
where #[a,b]\#[a,b] is the length of $[a,b]$, which is $\frac{pi}{2}$ for equation (26), and the number of samples (random points) we draw from a uniform distribution $\mathcal{U}(a,b)=\mathcal{U}(0,\frac{\pi}{2})$ is $k=50$. So for every point $j\in k$, we draw a random sample, we pass it to the neural network $\mathcal{F}$ to obtain the $\hat{y}$ and then we plug that in equation (28) to get the integral term $\mathcal{I}$. After we evaluate the integral term, we proceed in minimizing the loss function (27). The code listing 10 shows the Pytorch implementation of the loss function for this problem, where in lines 88-$12$, we evaluate the integral part described by Equation (28), and in line $15$, we compute the solution $\hat{y}$ using the DGM-like neural network. Finally, in line $17$, we estimate the loss function based on Equation (27).
Monte Carlo Integration It is a numerical method for estimating multidimensional integrals such as $\mathcal{I}=\int_{\Omega}f({\bf x})d{\bf x}$, where $\Omega\subseteq\mathbb{R}^{d}$. The naive algorithm for the Monte Carlo integration consists of the following steps: 1. Sample $k$ vectors ${\bf x}_{1},{\bf x}_{2},\ldots,{\bf x}_{k}$ from a uniform distribution on $\Omega$, ${\bf x}_{j}\sim\mathcal{U}(\Omega)$, 2. Estimate the term $\frac{\#\Omega}{k}\sum_{j=1}^{k}f({\bf x}_{j})$, where $\Omega=\int_{\Omega}d{\bf x}$. Due to the law of large numbers the sum eventually will converge to $\mathcal{I}$.
@@
We use the Adam optimizer with its default parameters and learning rate $\eta=0.0001$. We train the neural network for $3,000$ iterations. In each iteration we uniformly draw $n=32$ (batch size) random numbers from $0,\frac{\pi}{2}]$ representing the independent variable $x$ and we feed the mini-batch to the neural network.
@@
Figure 12 A shows the analytic solution of equation (26) (blue line) and the approximated solution evaluated with the trained neural network on the interval $0,\frac{\pi}{2}]$ with $50$ grid points. We see that the approximated solution overlays with the analytic one, and the MAE=$0.0134$. Panel 12 B shows the loss during training. We observe how the loss drops until convergence after about $500$ iterations.